# Benchmarking Vision Foundation Models for Input Monitoring in Autonomous Driving

### v2

@@ 4 Experiments
Datasets
Our experimental framework employs a diverse array of datasets to rigorously evaluate the robustness of our proposed method under various shifts. We carefully selected and curated datasets to simulate semantic shifts, covariate shifts, and their combination, providing a comprehensive testbed for our analysis.
@@
Pre-trained VFMs and Implementation Details
We employ a diverse selection of pre-trained VFMs, including CLIP [59], DINO [8], DINOv2 [54], and Grounding DINO [44], each representing a distinct paradigm in visual representation learning. CLIP [59] is evaluated across multiple architectures. Specifically, we utilize ResNet (RN) variants with varying depths (RN50, RN101) and width multipliers (RN50x4, RN50x16, RN50x64), alongside Vision Transformer (ViT) configurations with different patch sizes and model capacities (ViT-B/32, ViT-B/16, ViT-L/14) [^1] CLIP: https://github.com/openai/CLIP. DINO is examined through a range of backbones, including ViT variants with diverse patch sizes (Dino-ViT-S/16, Dino-ViT-S/8, Dino-ViT-B/16, Dino-ViT-B/8) and a ResNet variant (Dino-RN50) [^2] DINO: https://github.com/facebookresearch/dino. DINOv2 [54] is represented by Vision Transformer variants of increasing model capacity (Dinov2-ViT-S/14, Dinov2-ViT-B/14, Dinov2-ViT-L/14, Dinov2-ViT-G/14) [^3] DINOv2: https://github.com/facebookresearch/dinov2. Grounding DINO [44] employs a Swin Transformer [47] as its image encoder. In our experiments, we include Grounding DINO to evaluate the effectiveness of the Swin architecture [47] in comparison to ViT-based models, providing a broader perspective on the performance of different backbone architectures [^4] Grounding DINO: https://github.com/IDEA-Research/GroundingDINO.
As a baseline, we compare the performance of the VFMs against their ImageNet-trained counterparts [17]. This baseline includes ResNet variants (RN50, RN101, RN152) [25], ViT architectures (ViT-B/16, ViT-L/16) [19], and Swin Transformer backbones (SwinB, SwinT) [^5] ImageNet-trained ResNets: https://github.com/pytorch/pytorch; ViT and Swin Transformers: https://github.com/huggingface/pytorch-image-models [47]. Furthermore, we trained autoencoder models (AE) on the primary $\mathcal{X}_{\mathrm{monitor}}$ dataset to serve as feature extractors, offering an additional point of comparison with the VFMs. Detailed architectural specifications and implementation considerations are presented in the Appendix.
Semantic Shift
@@
APS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50x64 1024 0.94 0.93 0.24 0.94 0.93 0.24 0.94 0.93 0.23 0.88 0.84 0.37
ViT-B/16 512 0.88 0.86 0.38 0.89 0.88 0.43 0.91 0.89 0.33 0.91 0.88 0.34
DINO ViT-B/16 768 0.79 0.75 0.52 0.79 0.76 0.59 0.84 0.82 0.55 0.75 0.71 0.76
ViT-B/8 768 0.81 0.76 0.49 0.77 0.75 0.56 0.87 0.86 0.45 0.82 0.81 0.59
DINOv2 ViT-L/14 1024 0.89 0.87 0.43 0.84 0.82 0.68 0.82 0.77 0.52 0.83 0.78 0.54
ViT-G/14 1536 0.90 0.87 0.30 0.85 0.82 0.63 0.86 0.78 0.39 0.83 0.75 0.46
Grounding DINO SwinB 768 0.88 0.84 0.47 0.84 0.80 0.54 0.91 0.89 0.32 0.95 0.94 0.16
SwinT 768 0.85 0.80 0.53 0.83 0.79 0.59 0.92 0.90 0.28 0.96 0.95 0.16
@@
CLIP RN50x64 1024 0.93 0.94 0.54 0.93 0.95 0.54 0.94 0.95 0.53 0.94 0.96 0.43
ViT-B/16 512 0.91 0.93 0.50 0.91 0.93 0.54 0.92 0.94 0.48 0.93 0.95 0.49
DINO ViT-B/16 768 0.81 0.82 0.69 0.83 0.85 0.75 0.86 0.88 0.60 0.84 0.87 0.80
ViT-B/8 768 0.82 0.84 0.64 0.83 0.86 0.70 0.86 0.89 0.59 0.88 0.91 0.65
DINOv2 ViT-L/14 1024 0.78 0.79 0.77 0.78 0.81 0.80 0.82 0.85 0.74 0.86 0.88 0.63
ViT-G/14 1536 0.79 0.80 0.73 0.81 0.83 0.69 0.82 0.85 0.71 0.86 0.89 0.69
Grounding DINO SwinB 768 0.60 0.56 0.83 0.58 0.54 0.85 0.76 0.68 0.58 0.99 0.99 0.02
SwinT 768 0.57 0.54 0.89 0.57 0.54 0.91 0.64 0.58 0.78 0.91 0.84 0.20
Table 1: Comparison of VFMs with various backbones on the Lost and Found [57] and Bravo-Synobj[49] datasets, focusing on semantic shift sensitivity. Results show AUROC↑, AUPR↑ (higher is better) and FPR95↓ (lower is better) metrics. Best values per metric in bold.
@@
Real vs. Synthetic Data: Analysis of FPR95 scores demonstrates the effectiveness of Grounding DINO with Swin Transformer when combined with normalizing flows for distribution modeling. This architecture-method combination achieves notably low FPR across both real-world and synthetic datasets (Lost and Found: FPR95: 0.16; Bravo-Synobj: FPR95: 0.02), outperforming conventional approaches. While CLIP’s RN50x64 shows consistent performance with density estimation methods like GMM and OC-SVM (FPR95: 0.23–0.24 on Lost and Found), it exhibits higher false positive rates on synthetic data (FPR95: 0.53–0.54). These results suggest that Grounding DINO combined with flow-based density estimation provides effective modeling of feature distributions for semantic shift detection in both real and synthetic scenarios. Notably, simpler density estimation techniques (APS, OC-SVM) show limited effectiveness with Swin Transformer (FPR95: 0.47–0.54 on Lost and Found), highlighting the importance of selecting appropriate distribution modeling approaches for different architectural designs.
Architecture and Distribution Modeling: Our analysis reveals that Grounding DINO with Swin Transformer achieves exceptional performance when paired with flow-based density estimation, demonstrated by consistently low false positive rates across both real-world and synthetic scenarios (Lost and Found: FPR95: 0.16; Bravo-Synobj: FPR95: 0.02). The hierarchical design of Swin Transformer, with its local-global attention mechanism, appears to capture richer semantic relationships compared to standard ViT architectures. However, leveraging this enhanced semantic understanding requires appropriate distribution modeling – while CLIP’s RN50x64 maintains steady performance with conventional methods like GMM (FPR95: 0.23–0.24), Grounding DINO’s more nuanced representations necessitate flow-based density estimation, as evidenced by the limited effectiveness of simpler approaches (APS: FPR95: 0.47–0.54). This architecture-specific behavior suggests that the combination of Swin Transformer’s hierarchical feature learning and flow-based distribution modeling is particularly effective for semantic shift detection.
Model Complexity vs. Latent Dimensions: Notably, the strong detection capabilities of Grounding-DINO are achieved with a relatively compact latent space (768-dim) compared to CLIP’s RN50x64 (1024-dim), suggesting that architectural sophistication, rather than latent dimensionality, is the key driver of performance. The hierarchical feature learning in Swin Transformer appears to create more informative representations even in a lower-dimensional space, though these require more advanced density estimation techniques to be effectively leveraged.
Covariate Shift
Results for the two top-FPR5-scored backbones per model and superior GMM and NF methods are shown in Tab. 2 (find a full comparison of all combinations in the Appendix). The following key observations can be drawn from our experiments.
@@
VFMs demonstrate strong robustness to both semantic and covariate shifts, with most models achieving near-perfect FPR95 scores on both the SMIYC and Indian Driving Dataset. This is to be expected as both shifts at a time constitute a stronger deviation from the training data than previously. Hence, VFMs effectively handle complex real-world scenarios involving both novel objects and varying environmental conditions simultaneously.
Comparison with OOD Detection Methods
We evaluated our approach against established OOD detection methods [3] in the context of input monitoring for AD. To facilitate comparison, we implemented these methods using a DeepLabV3+ semantic segmentation network with ResNet50 backbone [^6] DeepLabv3+-RN50 trained on Cityscapes:: https://github.com/open-mmlab/mmsegmentation/tree/main/configs/deeplabv3plus trained on Cityscapes [15] dataset. Our VFM-based input monitoring technique was benchmarked against methods categorized by their detection mechanisms: feature-based approaches (GradNorm [33]) analyzing network activations and gradients; logit-based methods (Energy [45], Entropy [50], MaxLogit [88]) examining output characteristics; and distance-based approaches (KNN [77], DICE [76]) assessing feature space similarities, while ODIN [42] employs temperature scaling and input preprocessing. These methods require specific network access: intermediate layers for feature-based, final layer for logit-based, and feature embeddings for distance-based approaches. In contrast, our VFM-based method operates independently of the monitored network, significantly enhancing its applicability in AD systems where network architectures may be proprietary or inaccessible.
@@
Ours (RN50-GMM) 0.90 0.36 0.89 0.66 0.85 0.47 0.85 0.43
@@
Our comparative analysis reveals several insights about OOD detection methods across different types of distribution shifts. Traditional OOD detection methods show varying performance across semantic and covariate shifts. In particular, distance-based methods (DICE [76], KNN [77]) demonstrate moderate performance on Lost and Found (AUROC: 0.50-0.68) and Bravo-Synobj (AUROC: 0.49-0.57), suggesting their limited capability in capturing complex semantic shifts. This limitation likely stems from their reliance on direct feature space comparisons, which may not effectively capture high-level semantic anomalies.
Logit-based approaches (Energy [45], Entropy [50], MaxLogit [88]) perform consistently but modestly across all datasets (AUROC: 0.50-0.70), indicating their relative robustness but limited discriminative power. The uniform performance suggests these methods primarily capture model uncertainty rather than specific distribution shifts, making them less effective for targeted anomaly detection.
Notably, GradNorm [33] shows strong performance on covariate shifts (Foggy Cityscapes AUROC: 0.98, FPR95: 0.06; Bravo-Synrain AUROC: 0.99, FPR95: 0.00) and performs reasonably well on real semantic shifts (Lost and Found AUROC: 0.80, FPR95: 0.43) but struggles significantly with synthetic anomalies (Bravo-Synobj AUROC: 0.45, FPR95: 0.94). This contrasting performance suggests that gradient-based methods are particularly sensitive to global image transformations and can detect naturally occurring anomalies, but fail to identify synthetically inserted objects, likely due to their unnatural feature distributions that deviate from the gradient patterns observed in real-world scenes.
Input Monitoring
@@ 5 Conclusion
We presented an unsupervised, model-agnostic OOD detection framework that leverages latent representations of pre-trained off-the-shelf VFMs together with density modeling for joint detection of both semantic and covariate distributional shifts. This is for the first time thoroughly tested in the challenging domain of AD. Our extensive experiments show that simple techniques for modeling the VFM feature distributions are sufficient to detect all types of OOD shifts with astonishingly high accuracy. VFMs clearly outperform non-foundation model latent representations, and even proved competitive to classical OOD detectors specialized on one shift type. Interestingly, semantic richness plays a larger role in OOD performance than latent space dimensionality, but requires more flexible distribution modeling techniques. Lastly, despite our method being model-agnostic, downstream perception modules exhibit clear performance improvements when filtering out data that deviates too much from the training distribution. This promises that VMF latent space distributions can help to identify inputs with higher risk of inducing errors in CV DNNs, even in complex open-world domains.
@@ Appendix A Discussion of the Scope
  • A.1: Why input monitoring (instead of pixel-wise OOD segmentation)
@@
A.1 Input Monitoring vs. Pixel-Level OOD Segmentation
@@
Practically, OOD segmentation can complement OOD for more efficient anomaly treatment, but not replace it:
  • In case there is a shift in one part of the image, the probability that there are further (undetected) shifts elsewhere in the image is very high. Therefore, a reprocessing of the complete image using a safety mechanism is recommended. This requires an accurate prediction of whether there is any anomaly in the image. And despite considerable progress in pixel-level OOD segmentation [4, 23, 79, 63, 51, 10, 37, 61, 23, 86, 22, 46], these are still challenged by a high rate of false positives (illustrated in Fig. 3). This lies in the nature of problem formulation: Ensuring that there is no false positive among the tremendous amount of pixels in an image requires way more accurate predictions than a single image level OOD detection. This issue is illustrated in Fig. 3 via representative examples of Mask2Anomaly [61], a leading method from the SMIYC benchmark [9].

### v3

@@ 4 Experiments
Datasets:
Our experimental framework employs a diverse array of datasets to evaluate the robustness of our proposed method under various shifts. We carefully selected and curated datasets to simulate semantic shifts, covariate shifts, and their combination, providing a comprehensive testbed for our analysis.
@@
Pre-trained VFMs and Implementation Details:
We employ a diverse selection of pre-trained VFMs, including CLIP [61], DINO [10], DINOv2 [56], and Grounding DINO [45], each representing a distinct paradigm in visual representation learning. CLIP [61] is evaluated across multiple architectures. Specifically, we utilize ResNet (RN) variants with varying depths (RN50, RN101) and width multipliers (RN50x4, RN50x16, RN50x64), alongside Vision Transformer (ViT) configurations with different patch sizes and model capacities (ViT-B/32, ViT-B/16, ViT-L/14) [^1] CLIP: https://github.com/openai/CLIP. DINO is examined through a range of backbones, including ViT variants with diverse patch sizes (Dino-ViT-S/16, Dino-ViT-S/8, Dino-ViT-B/16, Dino-ViT-B/8) and a ResNet variant (Dino-RN50) [^2] DINO: https://github.com/facebookresearch/dino. DINOv2 [56] is represented by Vision Transformer variants of increasing model capacity (Dinov2-ViT-S/14, Dinov2-ViT-B/14, Dinov2-ViT-L/14, Dinov2-ViT-G/14) [^3] DINOv2: https://github.com/facebookresearch/dinov2. Grounding DINO [45] employs a Swin Transformer [48] as its image encoder, building upon the DETR-based detection architecture [86]. In our experiments, we include Grounding DINO to evaluate the effectiveness of the Swin architecture [48] in comparison to ViT-based models, providing a broader perspective on the performance of different backbone architectures [^4] Grounding DINO: https://github.com/IDEA-Research/GroundingDINO.
We also compare the performance of the VFMs against their ImageNet-trained counterparts [18]. This baseline includes ResNet variants (RN50, RN101, RN152) [26], ViT architectures (ViT-B/16, ViT-L/16) [20], and Swin Transformer backbones (SwinB, SwinT) [^5] ImageNet-trained ResNets: https://github.com/pytorch/pytorch; ViT and Swin Transformers: https://github.com/huggingface/pytorch-image-models [48]. Furthermore, we trained autoencoder (AE) on the primary $\mathcal{X}_{\mathrm{monitor}}$ dataset to serve as feature extractors, offering an additional point of comparison with the VFMs. Architectural specifications, comparisons with ImageNet-trained counterparts, and implementation considerations are presented in the Appendix.
4.1 Semantic Shift
@@
APS MFS OC-SVM GMM NF
AUROC AUPR FPR95 AUROC↑ AUPR↑ FPR95↓ AUROC AUPR FPR95 AUROC AUPR FPR95 AUROC AUPR FPR95
CLIP RN50x64 1024 0.94 0.93 0.24 0.93 0.93 0.27 0.94 0.93 0.24 0.94 0.93 0.23 0.88 0.84 0.37
ViT-B/16 512 0.88 0.86 0.38 0.86 0.85 0.41 0.89 0.88 0.43 0.91 0.89 0.33 0.91 0.88 0.34
DINO ViT-B/16 768 0.79 0.75 0.52 0.72 0.64 0.60 0.79 0.76 0.59 0.84 0.82 0.55 0.75 0.71 0.76
ViT-B/8 768 0.81 0.76 0.49 0.73 0.64 0.59 0.77 0.75 0.56 0.87 0.86 0.45 0.82 0.81 0.59
DINOv2 ViT-L/14 1024 0.89 0.87 0.43 0.91 0.88 0.37 0.84 0.82 0.68 0.82 0.77 0.52 0.83 0.78 0.54
ViT-G/14 1536 0.90 0.87 0.30 0.89 0.86 0.37 0.85 0.82 0.63 0.86 0.78 0.39 0.83 0.75 0.46
Grounding DINO SwinB 768 0.88 0.84 0.47 0.82 0.76 0.52 0.84 0.80 0.54 0.91 0.89 0.32 0.95 0.94 0.16
SwinT 768 0.85 0.80 0.53 0.79 0.73 0.60 0.83 0.79 0.59 0.92 0.90 0.28 0.96 0.95 0.16
@@
CLIP RN50x64 1024 0.93 0.94 0.54 0.93 0.94 0.54 0.93 0.95 0.54 0.94 0.95 0.53 0.94 0.96 0.43
ViT-B/16 512 0.91 0.93 0.50 0.91 0.93 0.50 0.91 0.93 0.54 0.92 0.94 0.48 0.93 0.95 0.49
DINO ViT-B/16 768 0.81 0.82 0.69 0.81 0.82 0.69 0.83 0.85 0.75 0.86 0.88 0.60 0.84 0.87 0.80
ViT-B/8 768 0.82 0.84 0.64 0.82 0.84 0.64 0.83 0.86 0.70 0.86 0.89 0.59 0.88 0.91 0.65
DINOv2 ViT-L/14 1024 0.78 0.79 0.77 0.78 0.79 0.77 0.78 0.81 0.80 0.82 0.85 0.74 0.86 0.88 0.63
ViT-G/14 1536 0.79 0.80 0.73 0.79 0.80 0.73 0.81 0.83 0.69 0.82 0.85 0.71 0.86 0.89 0.69
Grounding DINO SwinB 768 0.60 0.56 0.83 0.61 0.56 0.82 0.58 0.54 0.85 0.76 0.68 0.58 0.99 0.99 0.02
SwinT 768 0.57 0.54 0.89 0.58 0.54 0.90 0.57 0.54 0.91 0.64 0.58 0.78 0.91 0.84 0.20
Table 1: Comparison of VFMs with various backbones on the Lost and Found [59] and Bravo-Synobj[50] datasets. focusing on semantic shift sensitivity. Results show AUROC↑. AUPR↑ (higher is better) and FPR95↓ (lower is better) metrics. Best values per metric in bold.
@@
Architecture-Specific Distribution Modeling: Our analysis reveals that Grounding DINO with Swin Transformer achieves notable performance when paired with flow-based density estimation (Lost and Found: FPR95: 0.16; Bravo-Synobj: FPR95: 0.02). The hierarchical design of Swin Transformer, with its local-global attention mechanism, captures richer semantic relationships than standard architectures. However, leveraging this enhanced semantic understanding requires appropriate distribution modeling—while simpler approaches show limited effectiveness with Swin Transformer (APS, MFS, OC-SVM: FPR95: 0.47–0.54), normalizing flows unlock its full potential. This architecture-specific behavior demonstrates the importance of matching distribution modeling techniques to architectural characteristics.
Real vs. Synthetic Performance: Notably, CLIP’s RN50x64 maintains consistent performance with conventional methods like GMM and OC-SVM on real-world data (FPR95: 0.23–0.24 on Lost and Found), but struggles with synthetic anomalies (FPR95: 0.53–0.54 on Bravo-Synobj). In contrast, Grounding DINO with flow-based modeling maintains effectiveness across both real and synthetic scenarios, suggesting superior generalization to diverse semantic shift types regardless of data origin.
4.2 Covariate Shift
As previously, due to space limitations we present results for only the two top-performing backbones per model in terms of FPR95 scores, focusing on GMM and NF methods due to their superior detection capabilities and FPR95’s sensitivity in distinguishing architectural performance. A comprehensive comparison is available in the Appendix. The following key observations can be drawn from our experiments.
@@ 4.3 Semantic and Covariate Shift
VFMs demonstrate strong robustness to both semantic and covariate shifts, with most models achieving near-perfect FPR95 scores on both the SMIYC and Indian Driving Dataset. This is to be expected as both shifts at a time constitute a stronger deviation from the training data than previously.
4.4 Benchmarking Against OOD Methods and ImageNet-Trained Counterparts
We evaluated our approach against both established OOD detection methods [3] and ImageNet-trained counterparts of the same architectures in the context of input monitoring for AD systems. To facilitate comparison, we implemented these methods using a DeepLabV3+ semantic segmentation network with ResNet50 backbone [^6] DeepLabv3+-RN50 trained on Cityscapes:: https://github.com/open-mmlab/mmsegmentation/tree/main/configs/deeplabv3plus trained on Cityscapes [16] dataset. Our VFM-based input monitoring technique was benchmarked against methods categorized by their detection mechanisms: feature-based approaches (GradNorm [32]), which analyze network activations and gradients; logit-based methods (Energy [46], Entropy [51], MaxLogit [89]), which examine output characteristics; confidence-based methods (DICE [75]), which leverage confidence scores derived from the model; distance-based approaches (KNN [76]), which assess feature space similarities; and ODIN [42], which enhances OOD detection through temperature scaling and input preprocessing. These methods require specific network access: intermediate layers for feature-based, final layer for logit-based, and feature embeddings for distance-based approaches. In contrast, our VFM-based method operates independently of the monitored network, significantly enhancing its applicability in AD systems where network architectures may be proprietary or inaccessible.
@@ 4.4 Benchmarking Against OOD Methods and ImageNet-Trained Counterparts
Image-Net Trained Counterparts RN50-GMM 0.76 0.62 0.84 0.75 0.92 0.28 0.94 0.23
RN50-NF 0.75 0.63 0.87 0.74 0.93 0.25 0.97 0.12
SwinB-NF 0.86 0.44 0.91 0.53 0.92 0.35 0.99 0.05
VFMs Ours (RN50-GMM) 0.90 0.36 0.89 0.66 0.85 0.47 0.85 0.43
@@
OOD detection methods show varying performance across semantic and covariate shifts. In particular, confidence-based methods (DICE [75]) and distance-based methods (KNN [76]) demonstrate moderate performance on Lost and Found (AUROC: 0.50-0.68) and Bravo-Synobj (AUROC: 0.49-0.57), suggesting their limited capability in capturing complex semantic shifts. This limitation likely arises from their reliance on model confidence scores (DICE) and feature similarity measures (KNN), which are often unreliable under high semantic variations.
Logit-based approaches (Energy [46], Entropy [51], MaxLogit [89]) perform consistently but modestly across all datasets (AUROC: 0.50-0.70), indicating their relative robustness but limited discriminative power. Their uniform performance suggests that while these methods leverage logit magnitudes to separate ID and OOD samples, they may struggle with fine-grained anomalies that do not cause strong logit deviations.
Notably, feature-based methods like GradNorm [32] show strong performance on covariate shifts (Foggy Cityscapes AUROC: 0.98, FPR95: 0.06; Bravo-Synrain AUROC: 0.99, FPR95: 0.00) and perform reasonably well on real semantic shifts (Lost and Found AUROC: 0.80, FPR95: 0.43) but struggle significantly with synthetic anomalies (Bravo-Synobj AUROC: 0.45, FPR95: 0.94), likely because GradNorm is more sensitive to global image transformations and natural distortions, while synthetic object insertions do not trigger strong gradient deviations.
ImageNet-trained models perform well on covariate shifts but lag behind VFMs on semantic shifts. Using identical architectures (ResNet50, SwinB) trained solely on ImageNet with the same distribution modeling techniques, we found competitive performance on covariate shifts. SwinB-NF excelled on weather conditions (Bravo-Synrain: AUROC: 0.99, FPR95: 0.05) and RN50-NF performed strongly on fog detection (Foggy Cityscapes: AUROC: 0.93, FPR95: 0.25). However, a significant gap appears for semantic shifts, particularly with SwinB (ImageNet SwinB-NF on Bravo-Synobj: FPR95: 0.53 vs. VFM SwinB-NF: FPR95: 0.02). This suggests ImageNet pre-training captures features sufficient for environmental variations, while VFMs’ diverse training objectives and broader data exposure significantly enhance detection of novel objects and semantic anomalies.
4.5 Input Monitoring
@@ 5 Conclusion
We presented an unsupervised, model-agnostic binary OOD classification framework that leverages latent representations of pre-trained off-the-shelf VFMs together with density modeling for joint detection of both semantic and covariate shifts. This is for the first time thoroughly tested in the challenging domain of AD. Our extensive experiments show that simple techniques for modeling the VFM feature distributions are sufficient to detect semantic and/or covariate OOD shifts with astonishingly high accuracy. VFMs clearly outperform non-foundation model latent representations, and classical OOD methods. Lastly, despite our method being model-agnostic, down-stream perception modules exhibit clear performance improvements when filtering out data that deviates too much from the training distribution. This promises that VMF latent space distributions can help to identify inputs with higher risk of inducing errors in computer vision DNNs, even in complex open-world domains.
@@ Appendix A Discussion of the Scope
  • A.1: Why input monitoring (Comparison with pixel-wise OOD segmentation)
@@
A.1 Input Monitoring and Pixel-Level OOD Segmentation
@@
Practically, input monitoring complements pixel-wise OOD segmentation by enabling more efficient anomaly detection and response for the following reasons:
  • In case there is a shift in one part of the image, the probability that there are further (undetected) shifts elsewhere in the image is very high. Therefore, a reprocessing of the complete image using a safety mechanism is recommended. This requires an accurate prediction of whether there is any anomaly in the image. And despite considerable progress in pixel-level OOD segmentation [4, 24, 78, 65, 53, 12, 37, 63, 24, 87, 23, 47], these are still challenged by a high rate of false positives (illustrated in Fig. 2). This lies in the nature of problem formulation: Ensuring that there is no false positive among the tremendous amount of pixels in an image requires way more accurate predictions than a single image level OOD detection. This issue is illustrated in Fig. 2 via representative examples of Mask2Anomaly [63], a leading method from the SMIYC benchmark [11]. While pixel-wise OOD segmentation provides detailed information about environmental anomalies, its practical deployment in safety-critical applications remains challenged by high false positive rates.